# A mechanism to generate varying speed of light via Higgs-dilaton coupling:Theory and cosmological applications

### v3

@@ I Motivation
Against this backdrop, the accustomed Planck $cGh$ unit system (1899-1900) solidifies the role of the speed light $c$, the (Newton) gravitational constant $G$, and the quantum of action $\hbar$ as “fundamental constants” which encompass the realms of special relativity, gravitation, and quantum mechanics, respectively [7, 8, 9]. The trio $\{c,\,G,\,\hbar\}$ are viewed as fixed cornerstones, convertible to the three “fundamental units”—the Planck mass, the Planck length, and the Planck time, defined as
@@ VII Discussion and Summary
On scale symmetry—Symmetries plays a fundamental role in constructing modern physical theories, with notable examples including general covariance in GR and gauge invariance in the Glashow–Weinberg–Salam (GWS) model of particle physics. In his seminal 1995 report [155], Bardeen proposed employing scale symmetry as a viable alternative to supersymmetry to cure the naturalness problem of the Higgs mass. It is well established that the Higgs particles, as quanta of scalar fields, suffer a quadratic divergence in their loop amplitudes, which destabilizes their ultraviolet-limit behavior and requires fine tuning. Bardeen suggested that classical scale invariance could protect the Higgs mass from this divergence. Theories that impose scale invariance in the matter and gravity sectors have been actively pursued in both particle physics and gravitational physics [332, 298, 299, 300, 301, 274, 275, 276, 277, 228, 190, 189, 259, 247, 254, 327, 357, 353, 349, 258, 330, 162, 340, 333, 347, 348, 356, 311, 233, 260, 170, 307, 215, 238, 239, 249, 313, 303, 261, 278, 248, 335, 358, 337, 183, 309, 184, 186, 187, 185, 262, 263, 241, 328, 329, 235, 237, 234, 236, 346, 192, 191, 193, 194, 304, 217, 218, 219, 221, 222, 220, 314, 321, 176, 160, 202, 204, 207, 206, 201, 200, 203, 205, 280, 334, 316, 319, 326, 264, 265, 281, 282, 284, 283, 240, 250, 294, 230, 231, 285, 214, 352, 188, 227, 255, 266, 232, 286, 226, 224, 225, 315, 287, 310, 308, 267, 268, 177, 178, 180, 179, 181, 182, 288, 289, 290, 164, 169, 168, 167, 165, 166, 242, 243, 244, 245, 246, 223, 351, 251, 252, 253, 355, 197, 229, 302, 256, 171, 322, 323, 325, 324, 318, 345, 291, 216, 295, 163, 359, 211, 212, 213, 341, 279, 342, 343, 269, 336, 320, 270, 271, 312, 195, 196, 257, 317, 161, 172, 344, 199, 198, 272, 157, 156, 158, 159, 338, 339, 297, 354, 305, 306, 209, 210, 350, 292, 273, 331, 208, 173, 174, 175, 293, 296, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129]. Scale invariance typically necessitates (non-minimal) coupling of the Higgs field with gravity via a scalar field $\chi$—commonly referred to as a “dilaton”—which arises naturally in various theoretical contexts, such as Kaluza–Klein, string theory, and braneworld scenarios [12].
@@ VIII Conclusion
Lastly, as $\hbar$ and $c$ become variable in scale-invariant gravity, they can no longer serve as “fundamental units” within Planck’s $cGh$ system [7, 8, 9]. Appendix B introduces a new unit system suitable for variable $\hbar$ and $c$.

### v4

@@ I Motivation
Against this backdrop, the accustomed Planck $cGh$ unit system (1899-1900) solidifies the role of the speed light $c$, the (Newton) gravitational constant $G$, and the quantum of action $\hbar$ as “fundamental constants” which encompass the realms of special relativity, gravitation, and quantum mechanics, respectively [7, 8]. The trio $\{c,\,G,\,\hbar\}$ are viewed as fixed cornerstones, convertible to the three “fundamental units”—the Planck mass, the Planck length, and the Planck time, defined as
@@ VII Discussion and Summary
On scale symmetry—Symmetries plays a fundamental role in constructing modern physical theories, with notable examples including general covariance in GR and gauge invariance in the Glashow–Weinberg–Salam (GWS) model of particle physics. In his seminal 1995 report [154], Bardeen proposed employing scale symmetry as a viable alternative to supersymmetry to cure the naturalness problem of the Higgs mass. It is well established that the Higgs particles, as quanta of scalar fields, suffer a quadratic divergence in their loop amplitudes, which destabilizes their ultraviolet-limit behavior and requires fine tuning. Bardeen suggested that classical scale invariance could protect the Higgs mass from this divergence. Theories that impose scale invariance in the matter and gravity sectors have been actively pursued in both particle physics and gravitational physics [332, 299, 300, 301, 302, 275, 276, 277, 278, 228, 189, 188, 260, 247, 254, 327, 357, 353, 349, 258, 330, 161, 340, 333, 347, 348, 356, 311, 233, 261, 169, 308, 214, 238, 239, 249, 313, 304, 262, 279, 248, 335, 358, 337, 182, 183, 185, 186, 184, 263, 264, 241, 328, 329, 235, 237, 234, 236, 346, 191, 190, 192, 193, 305, 216, 217, 218, 220, 221, 219, 314, 321, 175, 159, 201, 203, 206, 205, 200, 199, 202, 204, 281, 334, 316, 319, 326, 265, 266, 282, 283, 285, 284, 240, 250, 295, 230, 231, 286, 213, 352, 187, 227, 255, 267, 232, 287, 226, 223, 224, 315, 288, 310, 309, 268, 269, 176, 177, 179, 178, 180, 181, 289, 290, 291, 163, 168, 167, 166, 164, 165, 242, 243, 244, 245, 246, 222, 351, 251, 252, 253, 355, 196, 229, 303, 256, 170, 322, 323, 325, 324, 318, 345, 292, 215, 296, 162, 359, 210, 211, 212, 341, 280, 342, 343, 270, 336, 320, 271, 272, 312, 194, 195, 257, 317, 160, 171, 344, 198, 197, 273, 156, 155, 157, 158, 338, 339, 298, 354, 306, 307, 208, 209, 350, 293, 274, 331, 207, 172, 173, 174, 294, 297, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 225, 259]. Scale invariance typically necessitates (non-minimal) coupling of the Higgs field with gravity via a scalar field $\chi$—commonly referred to as a “dilaton”—which arises naturally in various theoretical contexts, such as Kaluza–Klein, string theory, and braneworld scenarios [11].
@@ VIII Conclusion
Lastly, as $\hbar$ and $c$ become variable in scale-invariant gravity, they can no longer serve as “fundamental units” within Planck’s $cGh$ system [7, 8]. Appendix B introduces a new unit system suitable for variable $\hbar$ and $c$.